A bond with seven years to maturity pays a 5.5% coupon semiannually. The par value of the bond is $1,000 and the current price is $971.76. If you bought the bond today and held it to maturity you would earn 6%

If you sold it for $986.22 in one year, what would your total return be? Assume you collect both coupon payments.
Answer #1

Investor's total return comprises of 2 elements-capital gain/loss (change in market price) and coupon payment.

Rate of Return = (Sale value - cost of acquisituion + coupon pay) / cost of acquisituion

= (986.22-971.76+55)/971.76

= 69.46/971.76

= 7.15%

Total return in $ = Sale value - cost of acquisituion + coupon pay

= 986.22-971.76+55

= 69.46